# A Topic-wise Exploration of the Telegram Group-verse

### v1

@@ Abstract.
Although currently one of the most popular instant messaging apps worldwide, Telegram has been largely understudied in the past years. In this paper, we aim to address this gap by presenting an analysis of publicly accessible groups covering discussions encompassing different topics, as diverse as Education, Erotic, Politics, and Cryptocurrencies. We engineer and offer an open-source tool to automate the collection of messages from Telegram groups, a non-straightforward problem. We use it to collect more than 50 million messages from 669 groups. Here, we present a first-of-its-kind, per-topic analysis, contrasting the characteristics of the messages sent on the platform from different angles — the language, the presence of bots, the type and volume of shared media content. Our results confirm some anecdotal evidence, e.g., clues that Telegram is used to share possibly illicit content, and unveil some unexpected findings, e.g., the different sharing patterns of video and stickers in groups of different topics. While preliminary, we hope that our work paves the road for several avenues of future research on the understudied Telegram platform.
@@ 1. Introduction
Telegram has experienced remarkable growth in the past years, becoming one of the most popular instant messaging apps in the world. In July 2023, it surpassed the mark of 800 million monthly active users worldwide(tel, 2023). Telegram offers several features to its users, who can organize themselves into different spaces of communication such as groups (many-to-many) or channels (one-to-many).
Yet, the literature on Telegram is still limited in breadth. Targeting publicly accessible groups and channels, most prior works focused on textual content (e.g. news, hate speech), specific groups (e.g., terrorists (Khawaja, 2017)) or countries (e.g., Iran (Hashemi and Zare Chahooki, 2019)), and often a single topic of discussion (e.g., far-right politics (Urman and Katz, 2022)).
In contrast, we are driven by the hypothesis that user activity patterns on Telegram groups may be influenced by the main topic of discussion. We aim to offer a broad analysis of the platform usage by providing a first-of-its-kind, multi-faceted, topic-wise exploration of usage patterns on popular Telegram groups.
We use our crawler to gather data from more than a thousand open groups and select 669 of them with at least 100 active users, distributed across 10 different topics like Politics, Cryptocurrency, Video and Films, etc., where users discuss about some specific topic. In total our data covers around 51.6 M messages and 1.4 M distinct users over a two-month observation period.
We analyse our data aiming to precisely measure and contrast user activity patterns in groups across different topics, analysing the mix of various languages, the footprint of official Telegram bots, the diverse habits in sharing media (e.g., videos, audios, images, GIFs) and point to external content via URLs. We overall witness very different behavioural patterns; some are expected (e.g., the large usage of emojis in all topics and languages, or the sharing of possibly illicit content), others are more surprising (e.g., users in Darknet post much longer messages than users in the other topics; or videos published in Video and Films are much longer than those published in Erotic — the latter being with higher resolution).
Although preliminary, our results expose a very diverse universe. We believe that our work offers notable insights into the Telegram platform usage and constitutes a first step toward better understanding how users behave on such a platform.
@@ 2. Related work
The characteristics and dynamics of messaging platforms have attracted a lot of attention. Notably, prior studies analysed content properties (Resende et al., 2019b, a; Maros et al., 2020) and information spread (Caetano et al., 2019; Nobre et al., 2021) on WhatsApp’s groups, hinting at the catalytic role of the platform in various real-world events (Garimella et al., 2018; Bowles et al., 2020; Machado et al., 2019).
More recently, attention has been dedicated to groups and channels on Telegram, as the platform’s popularity increases across the globe. Some studies were interested in the inner workings of the mobile application (Satrya et al., 2016; Anglano et al., 2017), and its use by particular user populations, such as Iranian immigrants (Nikkhah et al., 2018), terrorist organizations (Khawaja, 2017), extremist groups (Hoseini et al., 2023), or particular countries (e.g., Iran and Russia (Hashemi and Zare Chahooki, 2019; Akbari and Gabdulhakov, 2019)). Others analysed the formation of communities within Telegram channels (Urman and Katz, 2022; Tikhomirova and Makarov, 2021) and their connection to information spread (Venâncio et al., 2024; Dargahi Nobari et al., 2021). Some other efforts studied content properties, limiting to textual content, and usage patterns, with attention given to news content (Naseri and Zamani, 2019), hate speech and abusive language (Wich et al., 2022), as well as the presence of fake channels (i.e., those impersonating important services or persons) (Morgia et al., 2023). The use of Telegram to perform illicit activities (e.g., pump-and-dump activities in cryptocurrency markets (Xu and Livshits, 2019), manipulation of social media popularity (Weerasinghe et al., 2020)) has also been previously addressed. Overall, previous works focused on the information people exchange on Telegram, in groups and channels of a specific topic. Only Morgia et al. (Morgia et al., 2023) used TGStat to gather channels associated with multiple topics. Yet, they did not distinguish between such topics and aggregated all of them to discover fake channels.
In contrast, we here offer a topic-wise analysis of common user activities in Telegram groups. Rather than focusing on the type of exchanged information, or how it spreads, we show how differently people leverage different features (e.g., media types, links to external sites) to interact with each other when discussing different topics.
@@ 3. Crawler and data collection
Given our interest in user interactions, our data collection effort is focused on Telegram public groups, i.e., public chats where all the members can send messages.
To collect the data, we design an open-source, two-stage crawler that we offer to the community. [^1] The code and data will be made available upon publication of the paper. At the first stage, the tool periodically crawls the TGStat website to discover public Telegram groups on various topics. At the second stage, the tool crawls Telegram, by joining the discovered groups and collecting all shared messages.
@@ 3.1. TGStat crawling
TGStat is a mostly undocumented service that catalogues popular Telegram groups and channels worldwide. Currently, TGStat’s database covers almost 1.9 M channels and groups (tgs, 2024), which are categorised into 48 pre-defined topics. For each topic, TGStat shows the lists of the top-100 groups/channels according to various metrics. As the whole database, these lists are dynamically updated. Each group is characterised by some metadata, including the group name, its topic, its language, and the monthly Active Users (AU), i.e., the number of unique users who have written at least one message inside the group in the past month.
We extract information from TGStat engineering a Python-based crawler using the BeautifulSoup Python package (bea, 2024). We periodically run the crawler to automatically extract the lists of groups. This allows us to grow the group lists in those topics of interest to us (see discussion in Section 4.1).
Most prior studies of Telegram searched for links to existing groups in social media, news and even word-of-mouth (Hoseini et al., 2023; Baumgartner et al., 2020; Dargahi Nobari et al., 2021). Such approach potentially leads to strenuous crawling work, especially if one is interested in covering diverse topics (as we do here). TGStat offers on a single platform easy access to a large number of already categorised groups, with dynamically updated activity metrics and topics. Such information allows us to narrow our focus to the most relevant groups, besides enabling a per-topic analysis. Indeed, a few recent studies have used TGStat as a starting point to explore Telegram (Tikhomirova and Makarov, 2021; Morgia et al., 2023; Urman and Katz, 2022). Yet, unlike these studies which gathered a single snapshot of listed groups, we here continuously monitor TGStat by growing the initial list of groups over multiple days. Moreover, given our interest in the per-topic analysis, we first estimate how reliable TGStat categorisation is, a neglected step in past works (see Section 4.1.2).
@@ 3.2. Telegram crawling
Given a list of previously discovered groups, our crawler automatises the group join and message collection tasks. We rely on the Telethon Python package (tel, 2024b) and design a scalable tool based on threads: a master instructs workers to join (and leave — if desired) a group, check if a pending request for join has been accepted, collect new messages, or just wait. For scalability, we use multiple Telegram IDs, each associated with a worker. The master maintains a list of groups to join and instruments workers to join these groups and collect all messages from a desired initial date until the present. We store the collected information in a MongoDB database for later processing.
We instrument our crawler to join and stay in groups discovered on TGStat. To refresh the collection of messages, workers download only the new messages since the last retrieved snapshot. For every group and message, the crawler stores all the returned information in JSON format in the MongoDB instance. In this paper, we focus on the following message information: sender user’s identifier, message body, message time, and possible media contained in the message (image, video, GIF, poll, etc.).
@@ 3.3. Crawler design challenges
Telegram implements several countermeasures to limit API abuse, notably: i) a limit of 500 groups a given Telegram ID can join; ii) an unspecified upper limit on the rate to join new groups which, if not respected, causes a lengthy temporary ban(tel, 2024a); iii) a without-any-notice permanent ban of novel-activated Telegram IDs. Respecting these limitations requires ingenuity when designing the crawler. First, we declared our intentions to the official Telegram support channel. Second, we carefully limited the group joining rate. Third, we used multiple Telegram IDs, each associated with a worker thread to scale the data gathering.
Telegram offers the possibility of setting up administration bots (known as Telegram bots) that ease group management. Captcha protection bots are popular for filtering fake user bots, i.e., actual Telegram accounts used to programmatically spam messages in open groups. Such captcha protection bots may kick users if they do not solve the captcha after a specific time. However, other bots or administrators might enforce different rules or criteria for group participation. Whenever we were removed from a group, we respected the administrators’ willingness and did not try to join the group again. Similarly, to respect the privacy indications of the group administrators, we only consider groups where “auto-delete” functionality is not enabled. [^2] The auto-delete setting erases messages for all participants either 24 hours or 7 days after sending.
Our crawler can collect up to thousands of messages per second per worker and join about tens of channels per hour without causing Telegram rate limitations.
@@ 4.1. Data collection and filtering
On April 1st, 2024 we collected the top-100 groups for all TGStat topics. Out of these, we select the subset of topics in which we are able to join at least 10 English language groups, a condition that allows us to manually validate the accuracy of TGStat’s topic labelling.
@@ 4.1.1. Topic selection
Out of the 48 topics, 12 topics have 10 or more English groups. We keep crawling TGStat every week to refresh the lists of top-100 groups for these 12 topics to observe how those lists change over time. We stop on May 1st, 2024, discovering 1,368 groups in total. We notice the largest growth in Erotics, Cryptocurrencies and Bookmaking, where we find around 20% new groups. This illustrates that taking a single snapshot from TGStat, as prior work (Tikhomirova and Makarov, 2021; Morgia et al., 2023; Urman and Katz, 2022), would limit the lists of discovered groups.
Feeding these growing lists to the Telegram crawler, 8.6% of groups have the auto-delete function enabled. We abandon them immediately. We instead fail to join 18.8% of groups because the group: (i) expired, or (ii) changed its name before we could join it, or (iii) is moderated and either the administrator did not admit us or a bot kicked us out after joining. We thus successfully join and monitor 993 groups overall. For each tracked group, we collect all messages starting from March 1st to April 30th. [^3] For this work, we limit the collection period to avoid overloading the Telegram servers. In total, we collect more than 50 M messages, with about 1 M new messages gathered each day.
Table 1. Dataset statistics for consistent topics. Column 2 refers to groups discovered in TGStat, all others refer to joined groups with at least 100 Active Users (AU).
Topic # Groups Avg ##\## Users ##\## Msg
@@
Total 1,154444The total number of groups adds up to 1,368 if we consider 214 groups from Courses and guides and Economics. 669 23.7 9.7 51.7 36.4
@@ 4.1.2. TGStat topic verification
As mentioned, we evaluate if the per-topic categorisation provided by TGstat is reliable. To that end, we check if the actual topic of discussion is (i) coherent with the topic assigned by TGStat and (ii) consistent across time. We pick all 206 English groups. For each group, we select three sets of 30 consecutive messages, each set separated by the others by ten days. We split the 206 groups into three partitions (with TGStat’s topics evenly distributed among partitions) and assign each to a human verifier who checks if the TGStat topic is consistent with the contents of the messages. In case of doubts, the verifier asks for the support of the other two.
TGStat’s topic assignment proves mostly correct, with two exceptions: the Courses and guides groups are mostly filled with spam; and the Economics groups mostly host discussions about cryptocurrencies, for which a dedicated topic already exists. We thus discard these two topics’ remaining 166 groups, ending up with 827 groups.
@@ 4.1.3. Selecting active groups
To guarantee groups are active and with enough diversity, we keep groups that have at least 100 active users, i.e. users who sent at least one message in the two-month observation period. From 827 groups, we discard 158 of them, ending with 669 groups, as detailed in column 3 of Table 4.1.1. For comparison, the total numbers of groups discovered in TGStat in each topic is shown in column 2. The table also details the average number of users per group, the percentage of active ones, the total number of messages and the average number of messages per active user. Figures vary widely, showing already very different interests, engagement and activity levels across the topics.
For the sake of completeness, Figure 7 in the Appendix B details the breakdown of the various cases one can face when trying to collect messages from Telegram. Depending on the topic, we observe various failure cases which may significantly reduce the number of groups to follow.

### v2

# Topic-wise Exploration of the Telegram Group-verse

@@ Abstract.
Although Telegram is currently one of the most popular instant messaging apps in the world, previous studies have mainly focused on analysing discussions on specific angles and topics. In this paper, we present a broad analysis of publicly accessible groups that cover a wide range of discussions, including Education, Erotic, Politics, and Cryptocurrencies. How do people interact with different topic groups? Is there any common or peculiar behaviour? We engineer and offer an open-source tool to automate the collection of messages from Telegram groups, a non-straightforward problem. We use it to collect more than 51 million messages from 669 groups. Here, we present a first-of-its-kind, per-topic analysis, contrasting the users’ activity patterns from different angles — the language, the presence of bots, the type and volume of shared media content, links to external platforms, etc. Our results confirm some anecdotal evidence, e.g., indications of spamming behaviour, and unveil some unexpected findings, e.g., the different sharing patterns of video and message length in groups of different topics. Our research provides a horizontal analysis of the public group in Telegram across various general topics, establishing a foundation for future studies that can delve deeper into user interactions and content dynamics within this unique messaging environment.
Telegram groups, Multimedia sharing, Topic characterization, User behavior
journalyear: 2025copyright: ccconference: Companion Proceedings of the ACM Web Conference 2025; April 28-May 2, 2025; Sydney, NSW, Australiabooktitle: Companion Proceedings of the ACM Web Conference 2025 (WWW Companion ’25), April 28-May 2, 2025, Sydney, NSW, Australiadoi: 10.1145/3701716.3717506isbn: 979-8-4007-1331-6/2025/04ccs: Information systems Social networksccs: Human-centered computing Social media
@@ 1. Introduction
Telegram has experienced remarkable growth in the past years, becoming one of the most popular instant messaging apps in the world. In July 2024, it surpassed the mark of 950 million monthly active users worldwide [^1] Telegram CEO Pavel Durov https://t.me/durov/337. Telegram offers several features to its users, who can organize themselves into different spaces of communication such as private chats (one-to-one), groups (many-to-many) or channels (one-to-many).
Yet, the literature on Telegram is still limited in breadth. Targeting publicly accessible groups and channels, most prior works focused on textual content (e.g. news, hate speech), specific groups (e.g., terrorists (Yayla and Speckhard, 2017)) or countries (e.g., Iran (Hashemi and Zare Chahooki, 2019)), and a single topic of discussion (e.g., far-right politics (Urman and Katz, 2022)).
In contrast, this paper is driven by the following research questions: (i) how do users behave on Telegram groups in terms of the features they most often make use of to interact with others (e.g., video sharing, link sharing, use of reactions, polls, etc.)? (ii) How do such platform usage patterns change across groups discussing different topics?
We develop an open-source crawler designed to access public Telegram groups, collect historical messages up to a specified date, and continuously update the message collection over time, to build a longitudinal archive. We use the crawler to gather data from more than a thousand open groups and focus on those with at least 100 active users, distributed across 10 different topics of discussion including Politics, Cryptocurrency, Darknet, Erotic, Video and Films, etc. In total our data covers around 51.6 M messages and 1.4 M distinct users over a two-month observation period.
We analyse our data aiming to measure and contrast user activity patterns in groups across different topics. We analyse the mix of various languages, the footprint of official Telegram bots, the diverse habits in sharing media (e.g., videos, audios, images, GIFs) and external content via URLs. We overall witness peculiar behavioural patterns; some might be expected (e.g., the presence of automated user behaviour to programmatically share context) while others are more surprising (e.g., users in Darknet post much longer messages than users in the other topics).
This paper is structured as follows. Section 2 lists related work. in Section 3, we present the architecture of our crawler. In Section 4, we detail the data collection process and present a first-of-its-kind characterisation of platform usage across different discussion topics. Section 5 further elaborates on the analysis of video content and URLs sharing patterns, together with a deeper analysis on the time elapsed between a YouTube video publication and its first appearance in Telegram groups under observation. Last, we draw conclusions in Section 6 and discuss ethics of the work in Section 7.
All in all, this work shows the heterogeneity of usages people exhibit on Telegram, corroborating known facts and exposing surprising findings. We hope this work stimulates other works in exploring some of the highlighted results. For this, we make both the data and the crawler open to the community.222 The code and data are available at https://anonymous.4open.science/r/TopicWiseTelegram-7A81
@@ 2. Related work
The characteristics and dynamics of messaging platforms have attracted a lot of attention. Notably, prior studies analysed content properties (Resende et al., 2019b, a; Maros et al., 2020) and information spread (Caetano et al., 2019; Nobre et al., 2022) on WhatsApp’s groups, hinting at the catalytic role of the platform in various real-world events (Garimella et al., 2018; Bowles et al., 2020; Machado et al., 2019).
More recently, attention has been dedicated to groups and channels on Telegram, as the platform’s popularity increases across the globe. Some studies were interested in the inner workings of the mobile application (Satrya et al., 2016; Anglano et al., 2017), and its use by particular user populations, such as Iranian immigrants (Nikkhah et al., 2018), terrorist organizations (Yayla and Speckhard, 2017; Alrhmoun et al., 2024), extremist groups (Hoseini et al., 2023; Kloo et al., 2024), or particular countries (e.g., Iran and Russia (Hashemi and Zare Chahooki, 2019; Akbari and Gabdulhakov, 2019; Hanley and Durumeric, 2024)). Others analysed the formation of communities within Telegram channels (Urman and Katz, 2022; Tikhomirova and Makarov, 2021) and their connection to information spread (Venâncio et al., 2024; Hoseini et al., 2024; Dargahi Nobari et al., 2021). Some other efforts studied content properties, limiting to textual content and usage patterns, with attention given to news content (Naseri and Zamani, 2019), hate speech and abusive language (Wich et al., 2022), as well as the presence of fake channels (i.e., those impersonating important services or persons) (Morgia et al., 2023). The use of Telegram to perform illicit activities (e.g., pump-and-dump activities in cryptocurrency markets (Xu and Livshits, 2019), manipulation of social media popularity (Weerasinghe et al., 2020)) has also been previously addressed. Overall, previous works focused on the information people exchange on Telegram, in groups and channels of a specific topic. Only Morgia et al. (Morgia et al., 2023) used TGStat to gather channels associated with multiple topics. Yet, they did not distinguish between such topics and aggregated all of them to discover fake channels.
In contrast, we here offer a topic-wise analysis of common user activities in Telegram groups. Rather than focusing on the type of information exchanged, or how it spreads, we show how people leverage different features (e.g., media types, links to external sites) to interact with each other, and how such patterns differ depending on the topics and goals of group discussion.
@@ 3. Crawler and data collection
Given our interest in user behaviour, we focus our data collection effort on Telegram public groups, i.e., public chats where all the members can send messages.
To collect the data, we design an open-source, two-stage crawler that we offer to the community.2 At the first stage, the tool periodically crawls the TGStat website to discover public Telegram groups on various topics. At the second stage, the tool crawls Telegram by joining the discovered groups and collecting all messages.
@@ 3.1. TGStat crawling
TGStat is a service that catalogues popular Telegram groups and channels worldwide. Currently, TGStat’s database covers almost 1.9 M channels and groups (tgs, 2024), which are categorised into 48 pre-defined topics. For each topic, TGStat shows the lists of the top-100 groups according to various metrics. These lists are continuously and dynamically updated. TGStat characterises each group by some metadata, including the group name, topic, language, and the monthly Active Users (AU), i.e., the number of unique users who have shared messages in the group in the past month.
We extract information from TGStat engineering a Python-based crawler using the BeautifulSoup package (bea, 2024). We periodically run it to automatically extract the lists of groups in various topics. This allows us to grow the group lists in those topics of interest to us (see discussion in Section 4.1).
Most prior studies of Telegram searched for links to existing groups in social media, news and even word-of-mouth (Hoseini et al., 2023; Baumgartner et al., 2020; Dargahi Nobari et al., 2021). This approach can demand extensive crawling, particularly for diverse topics. TGStat streamlines this process by offering categorized groups with real-time activity metrics, facilitating targeted selection and per-topic analysis. Prior studies (Tikhomirova and Makarov, 2021; Morgia et al., 2023; Urman and Katz, 2022) leveraged TGStat but relied on static snapshots. In contrast, we continuously expand our dataset over multiple days. Furthermore, given our focus on per-topic analysis, we assess the reliability of TGStat’s categorization—a crucial step overlooked in past work (see Section 4.1.2).
@@ 3.2. Telegram crawling
Given a list of previously discovered groups, our crawler automatises the group join and message collection tasks. We rely on the Telethon Python package (tel, 2024c) and design a scalable tool based on threads: a master instructs workers to join (and leave, if desired) a group, check if a pending request for join has been accepted, collect new messages, or just wait. For scalability, we use multiple Telegram IDs, each associated with a worker. The master keeps a list of groups to collect messages from and instructs workers to do so from a desired initial date until the present. We store the collected information in a MongoDB database for later processing.
We instrument our crawler to join and stay in groups as we discover them on TGStat. To refresh the collection of messages, workers download only the new messages since the last retrieved snapshot. For every group and message, the crawler stores all the returned information in JSON format in the MongoDB instance. In this paper, we focus on the following message information: sender user’s identifier, message body, message time, and media contained in the message (image, video, GIF, poll, etc.).
@@ 3.3. Crawler design challenges
Telegram implements several countermeasures to avoid API abuse, notably: i) a limit of 500 groups a given Telegram ID can join; ii) an unspecified upper limit on the rate to join new groups which, if not respected, causes a lengthy temporary ban (tel, 2024b); iii) a without-any-notice permanent ban of novel-activated Telegram IDs that start to interact with the platform with high frequency. Respecting these limitations requires ingenuity when designing the crawler. First, we declared our intentions to the official Telegram support channel. Second, we carefully controlled the group joining rate to limit the temporal ban. Third, we used multiple already-active Telegram IDs, each associated with a worker thread to scale the data gathering.
Telegram offers the possibility of setting up administration bots (known as Telegram bots) that ease group management. Captcha protection bots are popular for filtering fake user bots, i.e., actual Telegram accounts used to programmatically spam messages in open groups. Such captcha protection bots may kick users out if they do not solve the captcha after a specific time. Other bots or administrators might enforce different rules or criteria for group participation. Whenever we were removed from a group, we respected the administrators’ willingness and did not try to join the group again. Similarly, to respect the privacy indications of the group administrators, we only consider groups where the participants’ messages are persistent and not automatically removed (“auto-delete” functionality removes messages 24 hours or 7 days after sending).
Our crawler can collect up to thousands of messages per second per worker and join tens of groups per hour without overcoming Telegram rate limitations.
@@ 4.1.1. Topic selection
On April 1st, 2024 we collected the top-100 groups for all TGStat 48 topics. Out of these, we select the 12 topics in which we were able to join at least 10 English language groups, a condition that allows us to manually validate the accuracy of TGStat’s topic labelling. We keep crawling TGStat every week to refresh the lists of top-100 groups for these 12 topics to observe how those lists change over time. We stop on May 1st, 2024, discovering 1,368 groups in total. The growth in the number of groups discovered over time is notable in specific topics such as Erotics, Cryptocurrencies and Bookmaking, where we find around 20% new groups every week. This illustrates that taking a single snapshot from TGStat, as prior work (Tikhomirova and Makarov, 2021; Morgia et al., 2023; Urman and Katz, 2022), would limit the lists of discovered groups.
Feeding these growing lists to the Telegram crawler, we find that 8.6% of groups have the auto-delete function enabled. We abandon them immediately. We also fail to join 18.8% of groups because the group (i) did not exist anymore, (ii) changed their name before we could join it, or (iii) are moderated and either the administrator did not admit us or a bot kicked us out after joining. We thus successfully join and monitor 993 groups. For each tracked group, we collect all messages starting from March 1st to April 30th. [^3] For this work, we limit the collection period to avoid overloading the Telegram servers. In total, we collect over 50 M messages, with about 1 M new messages gathered each day.
Table 1. Dataset statistics for verified topics. Column 2 refers to groups discovered in TGStat, all others refer to joined groups with at least 100 Active Users (AU) with consistent topic.
Topic # Groups Average ##\## Users ##\## Messages
@@
Total 1,154444The total number of groups adds up to 1,368 if we consider 214 groups from Courses and guides and Economics that we initially considered and later discarded. 669 23.7 9.7 51.7 36.4
@@ 4.1.2. TGStat topic verification
Next, we verify if the per-topic categorisation provided by TGstat is reliable. To that end, we check if the actual topic of discussion is (i) coherent with the topic assigned by TGStat and (ii) consistent across time. We pick all English groups (206) and for each group, we select three sets of 30 consecutive messages, each set separated by the others by ten days. The 206 groups where randomly assigned to three human evaluators, each required to independently assess whether each set of messages was, collectively, addressing a subject that was consistent with the topic of the group. To measure the agreement among evaluators, we use the Fleiss’ Kappa (Fleiss et al., 1981). The labelling showed an agreement of 0.71, indicating substantial agreement between evaluators.
TGStat’s topic assignment proves mostly correct, with two exceptions: the Courses and guides groups are mostly filled with spam; and the Economics groups mostly host discussions about cryptocurrencies, for which a dedicated topic already exists. We thus discard the 166 groups in these two topics, ending up with 827 groups.
@@ 4.1.3. Selecting active groups
To guarantee groups are active and with enough diversity, we keep groups that have at least 100 active users, i.e. users who sent at least one message in the two-month observation period. From the 827 groups, we discard 158 of them, ending with 669 groups, as detailed in column 3 of Table 4.1.1. For comparison, the total number of groups discovered in TGStat in each topic is shown in column 2. The table also details the average number of users per group, the percentage of active ones, the total number of messages and the average number of messages per active user. Figures vary widely, showing already very different interests, engagement and activity levels across the topics.
For the sake of completeness, Figure 10 in the Appendix A details the breakdown of the various cases we faced when trying to collect messages from Telegram. Depending on the topic, we observe various failure cases which may significantly reduce the number of groups to follow.